# Make a Cheap Scaling: A Self-Cascade Diffusion Model for Higher-Resolution Adaptation

### v1

@@ 4.4 Analysis and Discussion
Drawing inspiration from previous explorations on scale adaptation [10], we found that directly applying the SD 2.1 model trained with $512^{2}$ images to generate $1024^{2}$ images led to issues such as object repetition and diminished composition capacity (see Figure 1). We observed that the local structural details of the generated images appeared reasonable and abundant without smoothness when the adapted scale was not large (e.g., 4×4$\times$4 × more pixels). In summary, the bottleneck for adapting to higher resolutions lies in the semantic component and composition capacity. Fortunately, the original pre-trained low-resolution diffusion model can generate a reliable low-resolution pivot, naturally providing proper semantic guidance by injecting the pivot semantic features during the higher-resolution diffusive sampling process. Simultaneously, the local structures can be completed based on the rich texture prior learned by the diffusion model itself, under strong semantic constraints.
Algorithm 1 Time-aware feature upsampler tuning.
@@
2:     $(z_{0}^{r},z_{0}^{r-1},c)\sim p(z^{r},z^{r-1},c)$
3:     $t\sim\text{Uniform}\{1,\ldots,K\}$
4:     $\epsilon\sim\mathcal{N}(\mathbf{0},\mathbf{I})$
5:     $z_{t}^{r}=\sqrt{\bar{\alpha}_{t}}z_{0}^{r}+\sqrt{1-\bar{\alpha}_{t}}\epsilon$
6:     $\theta_{\Phi}\leftarrow\theta_{\Phi}-\eta\bigtriangledown_{\theta_{\Phi}}\|\tilde{\epsilon}_{\theta+\theta_{\Phi}}(z_{t}^{r},t,c,z_{0}^{r-1})-\epsilon\|^{2}$
7:end while
8:$\textbf{return}\;\theta_{\Phi}$
Algorithm 2 Pivot-guided inference for $\mathbb{R}^{d_{r-1}}\rightarrow\mathbb{R}^{d_{r}}$.
@@
  • Minimal fine-tuning data. Previous cascaded model chains necessitate sequential, separate training, with each model being trained from scratch, thereby imposing a significant training burden. Our model is designed to quickly adapt the low-resolution synthesis model to higher resolutions using a small amount of high-quality data for fine-tuning.
  • Easy scaling. Our model possesses the extendable capacity for adapting to a newly requested higher resolution. This is achieved by re-tuning the integrated feature upsampling modules without necessitating any additional parameters. In contrast, previous cascading models require the training of an extra super-resolution model for such adaptations.
@@ 5 Experiments
Methods
# Trainable Param
Training Step
Infer Time
@@ Methods
[…]
Original
0
1×1$\times$1 ×
29.89
0.010
24.21
0.007
Attn-SF [16]
0
1×1$\times$1 ×
29.95
0.010
22.75
0.007
ScaleCrafter [10]
0
1×1$\times$1 ×
20.88
0.008
16.67
0.005
Ours-TF (Tuning-Free)
0
1.04×1.04$\times$1.04 ×
12.25 0.004 6.09 0.001
Full Fine-tuning ($18k$)
860M
$18k$
1×1$\times$1 ×
21.88
0.007
17.14
0.005
LORA-R32
15M
$18k$
1.22×1.$2$
17.02
0.005
11.33
0.003
LORA-R4
1.9M
$18k$
1.20×1.20$\times$1.20 ×
14.74
0.005
9.47
0.002
SD+SR
184M
1.25M
5×5$\times$5 ×
12.59
0.005
Ours-T (Tuning)
0.002M
$4k$
1.06×1.06$\times$1.06 ×
12.40 0.004 3.15 0.0005
[…]
5.1 Implementation Details
The proposed method is implemented using PyTorch and trained on two NVIDIA A100 GPUs. The original base diffusion model’s parameters are frozen, with the only trainable component being the integrated upsampling modules. The initial learning rate is $5\times 10^{-5}$. We used $1000$ diffusion steps $T$ for training, and $50$ steps for DDIM [26] inference. We set $N=4$ and $K=700$ for all experiments. We conduct evaluation experiments on text-to-image models, specifically Stable Diffusion (SD), focusing on two widely-used versions: SD 2.1 [6] and SD XL 1.0 [18], as they adapt to two unseen higher-resolution domains. For the original SD 2.1, which is trained with $512^{2}$ images, the inference resolutions are $1024^{2}$ and $2048^{2}$, corresponding to 4×4$\times$4 × and 16×16$\times$16 × more pixels than the training, respectively. We also conduct evaluation experiments on text-to-video models, where we select the LVDM [9] as the base model which is trained with $16\times 256^{2}$ videos ($16$ frames), the inference resolutions are $16\times 512^{2}$, 4×4$\times$4 × more pixels than the training. We have included the experiments for SD XL 1.0 in the supplementary material.
5.2 Evaluation on Image Generation
[…]
Comparison with State-of-the-Art. We conduct the comparison experiments on two settings, i.e., training-free and fine-tuning. For training-free methods, we compare our training-free version, denoted as Ours-TF, with the vanilla text-to-image diffusion model (Original) that directly samples the higher resolution images via the original checkpoint, as well as two training-free methods, i.e., Attn-SF [16] and ScaleCrafter [10]. Besides, we also compare our fine-tuning version, denoted as Ours-T, with the full fine-tuning model, LORA tuning, and Any-Size-Diffusion [37]. We fine-tune all competing methods by applying online filtering on Laion-5B for high-resolution images larger than the target. Table 5 and Table 2 show the quantitative results on Laion-5B [23] over 4×4$\times$4 × and 16×16$\times$16 × more pixels compared to base model. We show random samples from our method on adapted various resolutions in Figure 3. Visual comparisons with the competing methods are included in supplementary materiel.
[…]
Original 104.70 0.043 104.10 0.040
@@ 5.2 Evaluation on Image Generation
Ours-TF 38.99 0.015 34.73 0.013
Table 2: Quantitative results of different methods on the dataset of Laion-5B with 16×16$\times$16 × image scale adaptation on $2048^{2}$ resolution.
@@ 5.3 Evaluation on Video Generation
Figure 3: Visual examples of the higher-resolution adaptation to various higher resolutions, e.g., $1024^{2}$, $3072\times 1536$, $1536\times 3072$, and $2048^{2}$, with the pre-trained SD 2.1 trained with $512^{2}$ images, comparing to $1024^{2}$ results of Full fine-tuning (Full-FT) and LORA-R4 (right down corner). Please zoom in for more details.
@@
Table 3: Quantitative results of different methods on the dataset of Webvid-10M with 4×4$\times$4 × video scale adaptation on $16\times 512^{2}$ resolution (16 frames). $10k$ and $50k$ denote the training steps of each method.
[…]
Comparison with State-of-the-art. To comprehensively verify the effectiveness of our proposed method, we also conduct comparison experiments on a video generation base model [9]. To the best of our knowledge, this paper is the first to investigate the scale adaptation problem in video synthesis. Thus, we compare our method with a full fine-tuning model and LORA tuning (consisting of two different ranks, i.e., 4 and 32), as well as the previous tuning-free method, i.e., ScaleCrafter. Table 3 shows the quantitative results on Webvid-10M [1] and visual comparisons are shown in Figure 4. Our method achieves better FVD and KVD results in approximately $20\%$ of the training steps compared to the competing approaches. With the merits of the reuse of reliable semantic guidance from a well-trained low-resolution diffusion model, our method can achieve better object composition ability (e.g., the reaction between cat and yarn ball and the times square as shown in the second and fourth examples of Figure 4, respectively) and rich local structures compared to the competing methods (e.g., the details of the teddy bear as shown in the third example of Figure 4). In contrast, for full fine-tuning models, the issue of low saturation and over-smoothness requires many training steps to resolve and it is difficult to achieve results as good as those obtained with low-resolution models. Besides, the generated results of both full fine-tuning and LORA tuning methods will have motion shift or motion inconsistency issues as shown in the bag of astronaut in the first example of Figure 4, while our method can better maintain the original model’s temporal consistency capabilities, generating more coherent videos (video examples refer to supplementary material).
5.4 Network Analysis
Efficiency comparison. To demonstrate the training and sampling efficiency of our method, we compare our approach with selected competing methods in Table 5 for generating $1024^{2}$ resolution images on the Laion-5B dataset. Our model has only 0.002M trainable parameters, utilizing approximately the parameters compared to LORA-R4 (with a rank of 4). Although our proposed method requires a cascaded generation process, i.e., starting with low-resolution generation followed by progressively pivot-guided higher-resolution generation, the inference time of our method is similar to that of the original baseline (with a factor of 1.04×1.04$\times$1.04 × for the tuning-free version and 1.06×1.06$\times$1.06 × for the tuning version), resulting in virtually no additional sampling time. Besides, we present the FID and FVD scores for several methods every $5k$ iteration on image (Laion-5B) and video (Webvid-10M) datasets as shown in Figure 5. Our observations demonstrate that our method can rapidly adapt to the desired higher-resolution. By cyclically reusing the frozen diffusion base model and incorporating only lightweight upsampler modules, our approach maximally retains the generation capacity of the pretrained base model, resulting in improved fine-tuned performance.
Refer to caption Refer to caption
Figure 6: Visual quality comparisons between the training-free methods and ours on higher-resolution adaptation with $1024^{2}$ resolutions. Please zoom in for more details.
@@ 5.4 Network Analysis
Figure 7: Visual examples of video generation of the (a) low-resolution pivot samples generated by the pre-trained base model, (b) super-resolution result by SD-SR, and (c) high-resolution final output of our tuning approach. Please zoom in for more details.
@@
Tuning-free or fine-tuning? Although our tuning-free self-cascade diffusion model can inject the correct semantic information to higher-resolution adaptation, some extreme examples still make it difficult to completely suppress repetition issues and composition capabilities, such as repetitive legs and sofas as shown in Figure 6. Such failure case is particularly evident in the repetition of very fine-grain objects or texture, which is a common occurrence among all tuning-free competing methods, like Attn-SF [16] and ScaleCrafter [10]. By tuning plug-and-play and lightweight upsampler modules with a small amount of higher-resolution data, the diffusion model can learn the low-level details at a new scale.
@@
Relation to the super-resolution methods. We also compare our approach to using a pre-trained Stable Diffusion super-resolution (SD 2.1-upscaler-4$\times$) as post-processing, denoted as SD-SR, for the higher-resolution generation as shown in Table 5. Our approach achieves better performance and reduced inference time, even in a tuning-free manner (Ours-TF). In contrast, SD-SR still requires a large amount of high-resolution data for training a new diffusion model with around 184M extra parameters to be trained. Furthermore, our method not only increases the resolutions of pivot samples like SD-SR, but also explores the potential of the pre-trained diffusion model for fine-grained details generation and inheriting the composition capacity. We illustrate one example of video generation in Figure 7, where the low-resolution pivot sample from the base model predicts the object shift result across temporal frames; however, our approach can effectively correct such inconsistencies, which is not achievable by simply applying SD-SR.
@@
Limitations. Our proposed method can effectively adapt to a higher-resolution domain. However, it still has limitations. Since the number of parameters in the plug-and-play upsampler modules we insert is very small, there is an upper bound to the performance of our method when there is sufficient training data, especially when the scale gap is too large, e.g., higher resolution than $4k$ resolution data. We will further explore the trade-off between adaptation efficiency and generalization ability in future work.

### v2

@@ 4.3 Analysis and Discussion
Drawing inspiration from previous explorations on scale adaptation [12], we found that directly applying the SD 2.1 model trained with $512^{2}$ images to generate $1024^{2}$ images led to issues such as object repetition and diminished composition capacity (see Fig. 1). We observed that the local structural details of the generated images appeared reasonable and abundant without smoothness when the adapted scale was not large (e.g., 4×4$\times$4 × more pixels). In summary, the bottleneck for adapting to higher resolutions lies in the semantic component and composition capacity. Fortunately, the original pre-trained low-resolution diffusion model can generate a reliable low-resolution pivot, naturally providing proper semantic guidance by injecting the pivot semantic features during the higher-resolution diffusive sampling process. Simultaneously, the local structures can be completed based on the rich texture prior learned by the diffusion model itself, under strong semantic constraints.
Algorithm 1 Feature upsampler tuning process.
@@
2:    $(x_{0},c)\sim p(x_{0},c)$
3:    $z^{r}_{0}=E(x_{0})$
4:    $z^{r-1}_{0}=E(Downsample(x_{0}))$
5:    $t\sim\text{Uniform}\{1,\ldots,K\}$
6:    $\epsilon\sim\mathcal{N}(\mathbf{0},\mathbf{I})$
7:    $z_{t}^{r}=\sqrt{\bar{\alpha}_{t}}z_{0}^{r}+\sqrt{1-\bar{\alpha}_{t}}\epsilon$
8:    $\theta_{\Phi}\leftarrow\theta_{\Phi}-\eta\bigtriangledown_{\theta_{\Phi}}\|\tilde{\epsilon}_{\theta,\theta_{\Phi}}(z_{t}^{r},t,c,z_{0}^{r-1})-\epsilon\|^{2}$
9:end while
10:$\textbf{return}\;\theta_{\Phi}$
Algorithm 2 Inference process for $\mathbb{R}^{d_{r-1}}\rightarrow\mathbb{R}^{d_{r}}$.
@@
  • Less fine-tuning data. Previous cascaded model chains necessitate sequential, separate training, with each model being trained from scratch, thereby imposing a significant training burden. Our model is designed to quickly adapt the low-resolution synthesis model to higher resolutions using a small amount of high-quality data for fine-tuning.
@@ 5 Experiments
Table 1: Quantitative results of different methods on the dataset of Laion-5B with 4×4$\times$4 × adaptation on $1024^{2}$ resolution. The best results are highlighted in bold. Note that Ours-TF and Ours-T denote the tuning-free version and the upsampler tuning version, respectively. “#Param” denotes the number of trainable parameters and “Infer Time” denotes the inference time of different methods v.s. Direct Inference. pFIDr/pKIDr denote the patch- We put ‘-’ since FIDb/KIDb are unavailable for SD+SRfootnotemark: .
@@
  Direct Inference 0 - 1×1$\times$1 × 29.89 0.010 20.88 0.0070 24.21 0.007
Attn-SF [19] 0 - 1×1$\times$1 × 29.95 0.010 21.07 0.0072 22.75 0.007
ScaleCrafter [12] 0 - 1×1$\times$1 × 20.88 0.008 21.00 0.0071 16.67 0.005
Ours-TF (Tuning-Free) 0 - 1.04×1.04$\times$1.04 × 12.25 0.004 19.59 0.0071 6.09 0.001
Full Fine-tuning 860M $18k$ 1×1$\times$1 × 21.88 0.007 19.33 0.0077 17.14 0.005
LORA-R32 15M $18k$ 1.22×1.$2$ 17.02 0.005 18.65 0.0076 11.33 0.003
LORA-R4 1.9M $18k$ 1.20×1.20$\times$1.20 × 14.74 0.005 18.06 0.0074 9.47 0.002
SD+SR 184M 1.25M 5×5$\times$5 × 12.59 0.005 17.21 0.0053 - -
Ours-T (Tuning) 0.002M $4k$ 1.06×1.06$\times$1.06 × 12.40 0.004 15.35 0.0058 3.15 0.0005
footnotetext: We follow the same comparison settings of ScaleCrafter [12]. Since FIDb/KIDb are evaluated on the original low-resolution by down-sampling, the down-sampling results of SD+SR will be roughly the same as the reference real image set which denotes “zero distance”.
Table 2: Quantitative results of different methods on the dataset of Laion-5B with 16×16$\times$16 × image scale adaptation to $2048^{2}$ resolution. The best results are highlighted in bold. $10k$ and $20k$ denote the training steps used for tuning.
@@
Direct Inference 104.70 0.043 104.10 0.040
@@
Ours-TF (Tuning-Free) 38.99 0.015 34.73 0.013
Full Fine-tuning ($20k$) 43.55 0.014 41.58 0.012
LORA-R4 ($20k$) 50.72 0.020 51.99 0.019
Ours-T (Tuning) ($10k$) 18.46 0.005 8.99 0.001
5.1 Implementation Details
The proposed method is implemented using PyTorch and trained on two NVIDIA A100 GPUs. The original base diffusion model’s parameters are frozen, with the only trainable component being the integrated upsampling modules. The initial learning rate is $5\times 10^{-5}$. We used $1000$ diffusion steps $T$ for training, and $50$ steps for DDIM [29] inference. We set $N=4$ and $K=700$ for all experiments. We conduct evaluation experiments on text-to-image models, specifically Stable Diffusion (SD), focusing on two widely-used versions: SD 2.1 [7] and SD XL 1.0 [21], as they adapt to two unseen higher-resolution domains. For the original SD 2.1, which is trained with $512^{2}$ images, the inference resolutions are $1024^{2}$ and $2048^{2}$, corresponding to 4×4$\times$4 × and 16×16$\times$16 × more pixels than the training, respectively. We also conduct evaluation experiments on text-to-video models, where we select the LVDM [13] as the base model which is trained with $16\times 256^{2}$ videos ($16$ frames), the inference resolutions are $16\times 512^{2}$, 4×4$\times$4 × more pixels than the base resolution. We left the experiments for SD XL 1.0 in the supplementary.
5.2 Evaluation on Image Generation
Dataset and evaluation metrics. We select the Laion-5B [26] as the benchmark dataset which contains 5 billion image-caption pairs. We fine-tune all tuning-based competing methods by applying online filtering on Laion-5B for high-resolution images larger than the target resolution. We randomly sample $30k$ images with text prompts from the dataset and evaluate the generated image quality and diversity using the Inception Distance (FID) and Kernel Inception Distance (KID) metrics, which are measured between the generated images and real images, denoted as FIDr and KIDr. Following previous work [12], we sample $10k$ images when the inference resolution is higher than $1024^{2}$. Besides, to address the issue of squeezed resolutions in standard FIDr/KIDr, we randomly cropped local patches to calculate these metrics instead of resizing, denoted as pFIDr/pKIDr [4]. To ensure consistency in image pre-processing steps, we use the clean-fid implementation [20]. Since pre-trained models can combine different concepts that are not present in the training set, we also measure the FID and KID metrics between the generated samples under the base training resolution and inference resolution, denoted as FIDb and KIDb. This evaluation assesses how well our method preserves the model’s original ability when adapting to a new higher resolution.
Comparison with state-of-the-art. We conduct the comparison experiments on two settings, i.e., tuning-free and fine-tuning. For the tuning-free setting, we compare our tuning-free version, denoted as Ours-TF, with the vanilla text-to-image diffusion model (Direct Inference) that directly samples the higher resolution images via the original checkpoint, as well as two tuning-free methods, i.e., Attn-SF [19] and ScaleCrafter [12]. Besides, we also compare our fine-tuning version, denoted as Ours-T, with the full fine-tuning model, and LORA tuning (consisting of two different ranks, i.e., 4 and 32, denoted as LORA-R4 and LORA-R32). Tab. 1 and Tab. 2 show the quantitative results on Laion-5B [26] over 4×4$\times$4 × and 16×16$\times$16 × more pixels compared to base model. Our methods outperform existing methods in both tuning-free and fine-tuning settings, especially when adapting to a higher-resolution domain, e.g., 16×16$\times$16 × scale adaptation. Besides, with the merits of injecting newly acquired higher-resolution data for tuning, our tuning version can achieve a better and more robust generation performance, especially for 16×16$\times$16 × scale adaptation. We show random samples from our method (Ours-T) on adapted various resolutions in Fig. 3, comparing to results of Full fine-tuning (Full-FT) and LORA-R4. Our method can achieve not only high-quality image results with rich structural details and accurate object composition, e.g., the relationship between the bear and motorbike as shown in Fig. 3. Visual comparisons with the competing methods are included in the supplementary.
@@ 5.3 Evaluation on Video Generation
Figure 3: Visual examples of Ours-T (Tuning) on the higher-resolution adaptation to various higher resolutions, e.g., $1024^{2}$, $3072\times 1536$, $1536\times 3072$, and $2048^{2}$, with the pre-trained SD 2.1 trained with $512^{2}$ images, comparing to $1024^{2}$ results of Full fine-tuning (Full-FT) and LORA-R4 (right down corner: red dashed box). Please zoom in for more details.
Dataset and evaluation metrics. We select the Webvid-10M [1] as the benchmark dataset which contains 10M high-resolution collected videos. We randomly sample 2048 videos with text prompts from the dataset and evaluate the generated video quality and diversity using video counterpart Frechet Video Distance (FVD) [33] and Kernel Video Distance (KVD) [34], denoted as FVDr and KVDr.
Comparison with state-of-the-art. To comprehensively verify the effectiveness of our proposed method, we also conduct comparison experiments on a video generation base model [13]. Thus, we compare our method with a full fine-tuning model and LORA tuning with different ranks, as well as the previous tuning-free method, i.e., ScaleCrafter. Tab. 3 shows the quantitative results on Webvid-10M [1] and visual comparisons are shown in Fig. 4. Our method achieves better FVD and KVD results in approximately $20\%$ of the training steps compared to the competing approaches. With the merits of the reuse of reliable semantic guidance from a well-trained low-resolution diffusion model, our method can achieve better object composition ability (e.g., the reaction between cat and yarn ball and the times square as shown in the second and fourth examples of Fig. 4, respectively) and rich local structures compared to the competing methods (e.g., the details of the teddy bear as shown in the third example of Fig. 4). In contrast, for full fine-tuning models, the issue of low saturation and over-smoothness requires many training steps to resolve and it is difficult to achieve results as good as those obtained with low-resolution models. Besides, the generated results of both full fine-tuning and LORA tuning methods will have motion shift or motion inconsistency issues as shown in the bag of the astronaut in the first example of Fig. 4, while our method can better maintain the original model’s temporal consistency capabilities, generating more coherent videos (video examples refer to supplementary).
5.4 Network Analysis
Efficiency comparison. To demonstrate the training and sampling efficiency of our method, we compare our approach with selected competing methods in Tab. 1 for generating $1024^{2}$ resolution images on the Laion-5B dataset. Our model has only 0.002M trainable parameters, utilizing approximately the parameters compared to LORA-R4 (with a rank of 4). Although our proposed method requires a cascaded generation process, i.e., starting with low-resolution generation followed by progressively pivot-guided higher-resolution generation, the inference time of our method is similar to that of direct inference (with a factor of 1.04×1.04$\times$1.04 × for the tuning-free version and 1.06×1.06$\times$1.06 × for the tuning version), resulting in virtually no additional sampling time. Besides, we present the FID and FVD scores for several methods every $5k$ iteration on image (Laion-5B) and video (Webvid-10M) datasets as shown in Fig. 6. Our observations demonstrate that our method can rapidly adapt to the desired higher resolution. By cyclically reusing the frozen diffusion base model and incorporating only lightweight upsampler modules, our approach maximally retains the generation capacity of the pretrained base model, resulting in improved fine-tuned performance.
Refer to caption Refer to caption
Figure 6: Visual quality comparisons between the tuning-free methods and ours on higher-resolution adaptation to $1024^{2}$ resolutions. Please zoom-in to see more details.
Table 3: Quantitative results of different methods on the dataset of Webvid-10M [1] with 4×4$\times$4 × video scale adaptation on $16\times 512^{2}$ resolution (16 frames). The best results are highlighted in bold. $10k$ and $50k$ denote the training steps used for tuning.
@@ 5.4 Network Analysis
Figure 7: Visual examples of video generation of the (a) low-resolution pivot samples generated by the pre-trained base model, (b) super-resolution result by SD+SR, and (c) high-resolution final output of our tuning approach.
@@
Tuning-free or fine-tuning? Although our tuning-free self-cascade diffusion model can inject the correct semantic information to higher-resolution adaptation, some extreme examples still make it difficult to completely suppress repetition issues and composition capabilities, such as repetitive legs and sofas as shown in Fig. 6. Such failure case is particularly evident in the repetition of very fine-grain objects or texture, which is a common occurrence among all tuning-free competing methods, like Attn-SF [19] and ScaleCrafter [12]. By tuning plug-and-play and lightweight upsampler modules with a small amount of higher-resolution data, the diffusion model can learn the low-level details at a new scale.
@@
Relation to the super-resolution methods. We also compare our approach to using a pre-trained Stable Diffusion super-resolution (SD 2.1-upscaler-4$\times$) as post-processing, denoted as SD+SR, for the higher-resolution generation as shown in Tab. 1. Our approach achieves better performance and reduced inference time, even in a tuning-free manner (Ours-TF). In contrast, SD+SR still requires a large amount of high-resolution data for training a new diffusion model with around 184M extra parameters to be trained. Furthermore, our method not only increases the resolutions of pivot samples like SD+SR, but also explores the potential of the pre-trained diffusion model for fine-grained details generation and inheriting the composition capacity. We illustrate one example of video generation in Fig. 7, demonstrating two key advantages of our method over SD+SR: (1) while the low-resolution pivot sample from the base model predicts an “object shift” result across temporal frames, our method effectively corrects such inconsistencies, which is not achievable by simply applying SD+SR; (2) our approach excels in synthesizing finer details and textures compared to using SR solely as post-processing, as shown in the enhanced zoom-in on the tiger region.
@@
Limitations. Our method adapts well to higher-resolution domains but still has limitations. Since the number of parameters in the upsampler modules we insert is very small, there is an upper bound to the performance of our method when there is sufficient training data, especially when the scale gap is too large, e.g., higher resolution than $4k$ resolution data. We will further explore the trade-off between adaptation efficiency and generalization ability in future work.